OpenStudy (anonymous):

What are 3 perfect squares? Please help?

OpenStudy (yanasidlinskiy):

Perfect squares are is the same number that is multiplied twice...For example: 5*5=25 6*6=36 4*4=16

OpenStudy (anonymous):

Okay, thank you!!

OpenStudy (jessiegonzales):

The following is a list of perfect squares. 12 = 1 22 = 4 32 = 9 42 = 16 52 = 25 62 = 36 72 = 49 82 = 64 92 = 81 102 = 100 112 = 121 122 = 144 If a variable with an exponent has an even exponent then it is a perfect square. To get the square root, we simply divide the exponent by 2. For example x8 is a perfect square, its square root is x4 . x11 is not a perfect square. The following is a list of perfect squares. 12 = 1 22 = 4 32 = 9 42 = 16 52 = 25 62 = 36 72 = 49 82 = 64 92 = 81 102 = 100 112 = 121 122 = 144 If a variable with an exponent has an even exponent then it is a perfect square. To get the square root, we simply divide the exponent by 2. For example x8 is a perfect square, its square root is x4 . x11 is not a perfect square. The following is a list of perfect squares. 12 = 1 22 = 4 32 = 9 42 = 16 52 = 25 62 = 36 72 = 49 82 = 64 92 = 81 102 = 100 112 = 121 122 = 144 If a variable with an exponent has an even exponent then it is a perfect square. To get the square root, we simply divide the exponent by 2. For example x8 is a perfect square, its square root is x4 . x11 is not a perfect square.

OpenStudy (yanasidlinskiy):

Your WElcome~!!:)